OpenStudy (anonymous):

If (x + 2) is a factor of x^3 − 6x^2 + kx + 10, k=? Medal and fan!!

OpenStudy (campbell_st):

so if (x + 2) is a factor then x = -2 is a zero so substitute x = -2 into the equation... and find the value of k, that makes it zero.. hope it helps

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{ (x + 2) is a factor of f(x)= x^3 − 6x^2 + kx + 10} \hspace{.33em}\\~\\ \normalsize \text{ -2 must be the root of the function .}f(x) \hspace{.33em}\\~\\ f(-2)= 0 \hspace{.33em}\\~\\ f(-2)= (-2)^3 − 6(-2)^2 + k(-2) + 10 \hspace{.33em}\\~\\ \normalsize \text{so} \hspace{.33em}\\~\\ (-2)^3 − 6(-2)^2 + k(-2) + 10=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

(-2)^3=-8

OpenStudy (campbell_st):

\[ (-2)^3 -6 \times (-2)^2 + k \times (-2) + 10 = 0\] simplify it and then solve for k and (-2)^3 = -8

OpenStudy (anonymous):